The base of an isosceles triangle is 1 in. less than the length of one of the equal sides If the perimeter of the triangle is 20 inches find the length of each of the sides
Answer by mananth(16946)

equal side =x
base = x-1
..
sum of sides = 20
x+x+x-1=20
3x+1=20
-1 -1
3x=21
/3
x= 7 in.
...
7, 7, & 6 inches are the sides
...
m.ananth@hotmail.ca